• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:04:26 on August31,2009

    Entry number 287243

    keyword=Day shift summary

    Day Shift Summary:

    Shift workers: J. P. Chen (SL), W. Deconinck (TO), J. H. Lee (3rd).

    8:00 Recover from CHL problem. No beam, Hall in controlled access.
    Alexandre working on Compton.
    Kent and student(s) working on detector PMT noise issue.
    We (Walter and I) also took the opportunity to set back the camaras
    for the target temperature and fan frequency/JT valve.

    13:00 MCC called that they are ready to start tuning beam. At that time, only Kent and his company still working on detectors.
    13:30. Kent/co finished detector work. Hall Controlled Access ended.
    13:45 In Beam Permit.
    14:20 Beam tuning done. Moved from empty to LH2 target
    14:20-14:40 ramp beam from 0 uA to 50 uA. Took a run (11829) to check detector.
    Then ramp to 70 uA.
    14:40 another run 11830 to check detectors. People (experts) are changing HVs and chekcing. Beam off at 14:45.
    14:50 Will start Moller measurements and spin dance now.
    Moving target to Empty.
    14:50-16:00 setting up and starting Moller.